OpenStudy (anonymous):

Two fractions have denominators 9 and 3. Their sum is 5/9. If the numerators are switched, the sum is 7/9. Determine the two fractions

OpenStudy (anonymous):

so i got x/9 + y/3 = 5/9 and y/9 + x/3 = 7/9

OpenStudy (anonymous):

that's good... now solve the system.... clear the fraction in both equations by multiplying by 9....

OpenStudy (anonymous):